# Neurons: Emulating the Human Visual Cortex Improves Fidelity and Interpretability in fMRI-to-Video Reconstruction

### v2

@@ 1 Introduction
Extensive experiments demonstrate the superior performance of Neurons, which significantly outperforms state-of-the-art (SOTA) baselines, particularly in video-based metrics. Neurons achieves a 0.196 improvement (26.6%) in spatiotemporal consistency metrics and an average improvement of 0.042 (19.1%) in semantic-level metrics. The benefits of the proposed sub-tasks are evident, as they enable more accurate shape, location, and semantic information extraction. Moreover, by decoupling these sub-tasks, Neurons enhances the decoding of brain activity. We map the projections of different sub-tasks to the brain’s visual region map, revealing functional consistency with the human visual cortex.
@@ 7 Conclusion
In this study, we present Neurons, a novel framework inspired by the hierarchical organization of the human visual cortex, designed to enhance the fidelity and interpretability of fMRI-to-video reconstruction. By decomposing the learning process into four specialized tasks, we effectively emulate the functional specialization of the human visual system. This hierarchical approach enables the framework to capture a broad spectrum of video content, ranging from low-level visual features to high-level semantic concepts. Our comprehensive experimental results demonstrate that Neurons surpasses previous methods across both video-based and frame-based evaluation metrics. The framework achieves marked improvements in spatiotemporal consistency, semantic accuracy, and pixel-level reconstruction quality. Moreover, the decoupled task structure not only enhances reconstruction performance but also provides interpretable insights into the functional alignment between the model and the human visual cortex. By mapping the outputs of each task to corresponding brain regions, we observed a strong correlation with the hierarchical processing of visual information in the human brain. This alignment further highlights its potential for brain-computer interfaces and clinical applications.

### v3

@@ 1 Introduction
Extensive experiments show that Neurons outperforms state-of-the-art baselines, especially on video-based metrics. It improves spatiotemporal consistency by 0.196 (26.6%) and semantic metrics by 0.042 (19.1%) on average. The proposed sub-tasks enhance shape, location, and semantic extraction, and their decoupling boosts brain activity decoding. Mapping sub-task projections onto the brain’s visual regions reveals functional alignment with the human visual cortex.
@@ 7 Conclusion
We present Neurons, a novel framework inspired by the hierarchical organization of the human visual cortex, designed to enhance the fidelity and interpretability of fMRI-to-video reconstruction. By decomposing the learning process into four specialized tasks, we effectively emulate the functional specialization of the human visual system. This hierarchical approach enables the framework to capture a broad spectrum of video content, ranging from low-level visual features to high-level semantic concepts. Our comprehensive experimental results demonstrate that Neurons surpasses previous methods across both video-based and frame-based evaluation metrics. The framework achieves marked improvements in spatiotemporal consistency, semantic accuracy, and pixel-level reconstruction quality. The decoupled task structure not only enhances reconstruction performance but also provides interpretable insights into the functional alignment between the model and the human visual cortex. By mapping the outputs of each task to corresponding brain regions, we observed a strong correlation with the hierarchical processing of visual information in the human brain, which further highlights its potential for brain-computer interfaces and clinical applications.
Acknowledgements
This work was supported by a research grant from the Joint Research Scheme (JRS) under the National Natural Science Foundation of China (NSFC) and the Research Grants Council (RGC) of Hong Kong (Project No. N_HKUST654/24), as well as a grant from the Research Grants Council of the Hong Kong Special Administrative Region, China (Project No. R6005-24).